If you purchased a property for $42500  14 years ago, what is your expected annual rate of return if you can sell the property for $838000 today and you assume annually compounding?    (Express your answer to two decimal places such as xx.xx. Do not include the % sign.)
Answer #1

Rate of return is given as equal to=(Future Value/Present Value)^(1/t)-1=(838000/42500)^(1/14)-1=23.73%